### v2

# Dirichlet-to-Neumann operator for the Helmholtz problem with general wavenumbers on the $n$-sphere

Abstract
This paper considers the Helmholtz problem in the exterior of a ball with Dirichlet boundary conditions and radiation conditions imposed at infinity. The differential Helmholtz operator depends on the complex wavenumber with non-negative real part and is formulated for general spatial dimensions. We prove wavenumber explicit continuity estimates of the corresponding Dirichlet-to-Neumann ($\operatorname*{DtN}$) operator which do not deteriorate as the complex wavenumber tends to zero.
The exterior Helmholtz problem can be equivalently reformulated on a bounded domain with $\operatorname*{DtN}$ boundary conditions on the artificial boundary of a ball. We derive wavenumber independent trace and Friedrichs-type inequalities for the solution space in wavenumber-indexed norms.
Keywords: Helmholtz equation, exterior problem, Dirichlet-to-Neumann, Sommerfeld radiation
AMS Classification: 31B10, 35J05, 47G10
1 Introduction
Acoustic scattering problems modelled by the Helmholtz equation are fundamental for the understanding of acoustic wave propagation and the design of efficient numerical methods. In many real-world applications such as design of antennas, radar detection, ultrasound imaging, and seismic wave analysis, the computational domain is the unbounded exterior of a bounded scatterer or of an emitting source. This unboundedness presents additional challenges in both theoretical analysis and the design of effective solution methods.
A popular approach to address this difficulty is to introduce a sufficiently large ball that contains the compact, possibly, non-homogeneous material and the support of a compactly supported source term. The boundary of this ball serves as an artificial boundary to the original exterior problem, restricting the Helmholtz equation to its interior. The key idea is to impose transparent boundary conditions on the artificial boundary such that a solution to the original problem (restricted to the interior of the ball) coincides with the solution on the finite domain, cf. [5, 7, 9, 10, 12, 15, 18] to only name a few.
This approach relies on the well-posedness of the unbounded problem, which, for exterior problems, requires appropriate conditions towards infinity, such as the Sommerfeld radiation condition [16]. Under such conditions, the solution to the homogeneous Helmholtz problem in the exterior of the ball is uniquely determined by its Dirichlet data on the artificial boundary. The Dirichlet-to-Neumann (DtN) operator maps a given Dirichlet trace to the normal derivative of the corresponding unique solution at the artificial boundary. The transparent boundary condition — also called $\operatorname*{DtN}$ boundary condition — on the artificial boundary requires the Dirichlet and Neumann traces of the truncated solution to match those of the original solution, which are expressed by the DtN operator. In this way, the DtN operator allows one to trade a finite computational domain for an additional boundary condition.
The theoretical justification and practical implementation of this approach require a thorough understanding of the involved DtN operator. In particular, for the numerical discretization of scattering problems, it is crucial to understand the explicit behaviour with respect to the wavenumber of the arising operators. Although the Fredholm theory in combination with (variants of) the unique continuation principle is a potent tool to establish well-posedness of acoustic problems, wavenumber explicit estimates require more refined analytical techniques. Among them are subtle definiteness (sign) properties of the real and imaginary part of the DtN operator [7, 9, 14, 18]. However, these properties have only been established for purely imaginary wavenumbers in $n=2,3$ dimensions [2, 14, 15], which prevents a unified theory for general complex wavenumbers: Existing results on Helmholtz problem based on the DtN approach are typically restricted to purely imaginary wavenumbers for $n\in\{2,3\}$, see, e.g., [7, 18] for recent developments on general coefficient Helmholtz problems.
To close this gap, this paper presents a unified mathematical analysis of spherical DtN operators for Helmholtz problems with general wavenumbers in any spatial dimension and provides a clearer mathematical understanding of the important definiteness characteristics of the DtN operator. In particular, our results enable a unified analysis of generalised boundary layer operators for Helmholtz problems with general coefficients and wavenumbers for $n\geq 2$ dimensions in [9]. The new insights also facilitate an improved tuning of control parameters in numerical methods depending on the real and imaginary part of the wavenumber such as [1] in future research.
Main results and further contributions.
a) The wavenumber-dependent DtN operator admits an explicit Fourier series representation in terms of eigenfunctions of the Laplace-Beltrami operator and Bessel functions. A careful wavenumber-explicit control of the Fourier series coefficients enables sharp bounds on the real and imaginary part of the DtN operator for general wavenumbers and spatial dimensions. Our results extend the existing theory for purely imaginary wavenumbers, refining the bounds in [14, Sec. 3] for two dimensions and recovering the classical results [15, Eqn. (2.6.23)–(2.6.24)] for three dimensions.
b) The solution space of the Helmholtz equation in the truncated domain with transparent ($\operatorname*{DtN}$) boundary conditions on the artificial boundary satisfies a novel Friedrichs-type inequality in $n\geq 3$ dimensions. This enables wavenumber-explicit Sobolev and trace estimates in the natural weighted norms which do not deteriorate for small complex wavenumbers with non-negative real part. These estimates are the key to an explicit characterisation of the natural weighted trace norm for Helmholtz problems on exterior domains in [8, Thm. 3.4].
c) The enforced condition towards infinity affects the definition of the solution operator for the exterior Dirichlet Helmholtz problem, whose normal derivative defines the $\operatorname*{DtN}$ operator. We discuss various types of radiation conditions and derive the well-posedness of the Helmholtz problem in the exterior of a ball with Dirichlet conditions on its boundary. This extends the results developed in [13] as well as the well-posedness results for the $\operatorname*{DtN}$ operator in [3].
We expect that these results in the field of applied analysis can be useful for analysing numerical discretization methods for problems with transparent boundary conditions or approximations thereof.
Outline. Section 2 introduces the notation for the geometric setting and corresponding Sobolev spaces as well as some standard trace estimates.
Section 3 presents the exterior Helmholtz problem with Dirichlet boundary conditions and discusses the popular choices of the radiation conditions towards infinity. The well-posedness of the resulting exterior problems is established by Proposition 3.2. Its proof follows mainly the theory developed in [13] and is postponed to the appendix. Theorem 3.4 states the important definiteness properties of the DtN operator with sharp bounds on its real and imaginary parts. The Friedrichs-type inequality described in b) is formulated in Theorem 3.5 and leads in Corollary 3.6 to the announced wavenumber-explicit Sobolev and trace estimates.
Section 4 introduces series expansions of the $\operatorname*{DtN}$ operators for the different radiation conditions. Sharp wavenumber-explicit estimates of the series coefficients for non-zero wavenumbers are formulated as Theorem 4.1 and proved in Subsection 4.2. The special case of wavenumber zero (i.e., the Laplacian) requires extra care and is analysed in Subsection 4.3. This paper concludes with the proofs of Theorems 3.4 and 3.5 in Subsection 4.4.
2 Preliminaries
The set of natural numbers is denoted by :=$\mathbb{N}:=\mathopen{}\mathclose{{}\left\{1,2,\dots}\right\}$ and we write $\mathbb{N}_{0}:=\mathbb{N}\cup\mathopen{}\mathclose{{}\left\{0}\right\}$. The following subset of the complex numbers will be needed:
$\mathbb{C}_{>0}:=\mathopen{}\mathclose{{}\left\{z\in\mathbb{C}\mid\operatorname{Re}z>0}\right\},\quad\mathbb{C}_{\geq 0}:=\mathopen{}\mathclose{{}\left\{z\in\mathbb{C}\mid\operatorname{Re}z\geq 0}\right\},\quad\mathbb{C}^{\bullet}_{\geq 0}:=\mathbb{C}_{\geq 0}\backslash\mathopen{}\mathclose{{}\left\{0}\right\}.$
The physical problem is formulated in $\mathbb{R}^{n}$ for general dimension $n\in\mathopen{}\mathclose{{}\left\{2,3,\ldots}\right\}$. We consider Lipschitz domains $\omega\subset\mathbb{R}^{n}$ which are open and connected (not necessarily bounded) sets whose boundary is compact and locally the graph of a Lipschitz function [13, Def. 3.28]. The classical Lebesgue and Sobolev spaces on $\omega$ are denoted by $L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)$ and $H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ with norms $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}$ and $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)}$. For $\kappa\geq 0$, the space $H_{\operatorname*{loc}}^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ is given by all distributions $v$ on the compactly supported (relative to $\mathbb{R}^{n}$) and smooth functions $C_{\operatorname*{comp}}^{\infty}\mathopen{}\mathclose{{}\left(\omega}\right)$ such that $\varphi v\in H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ for all $\varphi\in C_{\operatorname*{comp}}^{\infty}\mathopen{}\mathclose{{}\left(\omega}\right)$. For the definition of Sobolev spaces $H^{\kappa}\mathopen{}\mathclose{{}\left(\partial\omega}\right)$ on $\mathopen{}\mathclose{{}\left(n-1}\right)$-dimensional boundaries $\partial\omega$ and their norm $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{H^{\kappa}\mathopen{}\mathclose{{}\left(\Gamma}\right)}$ we refer, e.g., to [13, pp. 96–99]. We define wavenumber-indexed norms [^1] For $s=0$, the quantity $\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right),0}=\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}$ is only a seminorm in $H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)$. for $s\in\mathbb{C}_{\geq 0}$ by
$\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right),s}$ $:=\sqrt{\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}+\mathopen{}\mathclose{{}\left|s}\right|^{2}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}}$ for all $\text{for all }v\in H^{1}\mathopen{}\mathclose{{}\left(\omega}\right),$ (2.1)
$\mathopen{}\mathclose{{}\left\|g}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(\partial\omega}\right),s}$ $:=\sqrt{\mathopen{}\mathclose{{}\left\|g}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(\partial\omega}\right)}^{2}+\mathopen{}\mathclose{{}\left|s}\right|\mathopen{}\mathclose{{}\left\|g}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\partial\omega}\right)}^{2}}$ for all $\text{for all }g\in H^{1/2}\mathopen{}\mathclose{{}\left(\partial\omega}\right).$ (2.2)
An immediate consequence is the estimate $\|\bullet\|_{H^{1/2}\mathopen{}\mathclose{{}\left(\partial\omega}\right)}\leq\|\bullet\|_{H^{1/2}\mathopen{}\mathclose{{}\left(\partial\omega}\right),s}$.
The space $H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right):=\mathopen{}\mathclose{{}\left\{u\in H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)\mid\Delta u\in L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}\right\}$ is equipped with the (semi-)norms
$\mathopen{}\mathclose{{}\left\|{v}}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right)}$ $:=\sqrt{\mathopen{}\mathclose{{}\left\|{v}}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}+\mathopen{}\mathclose{{}\left\|\Delta{v}}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}}$ for all $\text{for all }v\in H^{1}(\omega,\Delta),$
$\mathopen{}\mathclose{{}\left|{v}}\right|_{H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right)}$ $:=\sqrt{\mathopen{}\mathclose{{}\left\|\nabla{v}}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}+\mathopen{}\mathclose{{}\left\|\Delta{v}}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}}$ for all $\text{for all }v\in H^{1}(\omega,\Delta).$
It is well-known that the outer unit normal $n(x)\in\mathbb{S}_{1}$ of the Lipschitz domain $\omega\subset\mathbb{R}^{n}$ exists for almost every $x\in\partial\omega$. The classical normal derivative $\partial/\partial n:\overline{\omega}\rightarrow\Gamma$ on some relatively open connectivity component $\Gamma\subset\partial\omega$ of the boundary can be continuously extended to a continuous mapping $\gamma_{1}:H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right)\rightarrow H^{-1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)$ with
$\mathopen{}\mathclose{{}\left\|\gamma_{1}v}\right\|_{H^{-1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)}\leq C_{\omega,1}\mathopen{}\mathclose{{}\left|v}\right|_{H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right)}\quad\text{for all }v\in H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right).$ (2.3)
Also the following standard and multiplicative trace estimates are well known:
$\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)}\leq C_{\omega,0}\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)}\quad\text{and\quad}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Gamma}\right)}\leq C_{\omega,0}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{1/2}\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)}^{1/2}\quad\text{for all }v\in H^{1}\mathopen{}\mathclose{{}\left(\omega}\right).$ (2.4)
The space $H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right)$ is a subspace of
$H^{1}\mathopen{}\mathclose{{}\left(\omega,\Gamma}\right):=\mathopen{}\mathclose{{}\left\{v\in H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)\mid\gamma_{1}v\in H^{-1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)}\right\}$ (2.5)
which is equipped with the natural norm
$\mathopen{}\mathclose{{}\left\|\bullet}\right\|_{H^{1}(\omega,\Gamma)}:=\sqrt{\mathopen{}\mathclose{{}\left\|\bullet}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}+\mathopen{}\mathclose{{}\left\|\gamma_{1}\bullet}\right\|_{H^{-1/2}(\Gamma)}^{2}}.$
The space $L(H_{1};H_{2})$ of continuous linear maps $A:H_{1}\to H_{2}$ between Hilbert spaces $H_{1}$ and $H_{2}$ is equipped with the operator norm topology. For the Fréchet space $H^{1}_{\mathrm{loc}}(\omega)=H_{2}$ as the image space, we set
$L(H_{1};H^{1}_{\mathrm{loc}}(\omega)):=\{A:H_{1}\to H^{1}_{\mathrm{loc}}(\omega)\ |\ \varphi A\in L(H_{1};H^{1}(\omega))\text{ for all }\varphi\in C^{\infty}_{\mathrm{comp}}(\omega)\}.$
Let $B_{R}\subset\mathbb{R}^{n}$ denote the open ball of radius $R>0$ at the origin. The outer unit normal vector $x/\mathopen{}\mathclose{{}\left\|x}\right\|$ on the boundary $S_{R}=\partial B_{R}$ points into the unbounded complement $B_{R}^{+}:=\mathbb{R}^{n}\setminus\overline{B_{R}}$ of $B_{R}$ and its derivative in this direction is denoted by $\partial_{r}$. The notation ||$\bullet$|| ∙ | is context-sensitive and may refer to the Lebesgue measure $|\omega|$ of a bounded measurable $n$-dimensional set $\omega\subset\mathbb{R}^{n}$, the surface measure $|\Gamma|$ of an $\mathopen{}\mathclose{{}\left(n-1}\right)$-dimensional manifold $\Gamma\subset\mathbb{R}^{n}$, and the cardinality $|J|$ of a countable set $J$.
Remark 2.1.
Let $s\in\mathbb{C}_{\geq 0}$. Note that any function $u\in H_{\operatorname*{loc}}^{1}\mathopen{}\mathclose{{}\left(B_{R}^{+}}\right)$ with $\Delta u+s^{2}u=0$ in a bounded neighborhood $\omega$ of $S_{R}$ in $B_{R}^{+}$ satisfies $\mathopen{}\mathclose{{}\left.u}\right|_{\omega}\in H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right)$ and the normal derivative $\partial_{r}u$ of $u$ on $S_{R}$ is well defined in $H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$.
3 The Dirichlet-to-Neumann operator on a sphere
This section discusses the well-posedness of the Dirichlet Helmholtz problem on the exterior of the ball $B_{R}$ and the properties of the associated Dirichlet-to-Neumann operator (DtN) on the sphere $S_{R}$.
3.1 The Dirichlet Helmholtz problem on the exterior of a ball
Let $s\in\mathbb{C}_{\geq 0}$ denote the wavenumber and consider the exterior Dirichlet Helmholtz problem
find $\text{find }u\in H_{\operatorname*{loc}}^{1}\mathopen{}\mathclose{{}\left(B_{R}^{+}}\right)\quad\text{such that\quad}\mathopen{}\mathclose{{}\left\{\begin{array}[c]{rl}-\Delta u+s^{2}u=0&\text{in }B_{R}^{+}=\mathbb{R}\setminus\overline{B_{R}},\\ u=g&\text{on }S_{R}=\partial B_{R}.\end{array}}\right.$ (3.1)
These equations can be closed by prescribing the behaviour of $u$ by some conditions towards infinity (denoted by $\operatorname*{IC}_{s}$). These conditions are by no means unique; the first version is given by (3.2a), (3.2b) and denoted by $\operatorname*{IC}_{s}^{\operatorname*{strong}}$
$\mathopen{}\mathclose{{}\left\{\begin{array}[c]{rl}\mathopen{}\mathclose{{}\left|u\mathopen{}\mathclose{{}\left(r\xi}\right)}\right|&\leq Cr^{\mu}\\ \mathopen{}\mathclose{{}\left|\frac{\partial u}{\partial r}\mathopen{}\mathclose{{}\left(r\xi}\right)+su\mathopen{}\mathclose{{}\left(r\xi}\right)}\right|&\leq Cr^{\mu-1}\end{array}}\right\}\quad\text{as }r\rightarrow\infty\text{ uniformly in }\xi\in S_{1}.$ (3.2a)
The rate $\mu$ depends on the spatial dimension $n$ and the wavenumber $s:$
$\mu:=\mathopen{}\mathclose{{}\left\{\begin{array}[c]{ll}\frac{1-n}{2}&\text{if}s\in\mathbb{C}^{\bullet}_{\geq 0},\\ 2-n&\text{if }s=0.\end{array}}\right.$ (3.2b)
Remark 3.1 (infinity conditions).
There exist alternative conditions towards infinity in the literature [^2] The superscript “$\operatorname*{Green}$” indicates that for $\operatorname*{IC}_{s}^{\operatorname*{Green}}$ a Green’s representation theorem holds under certain assumptions while the superscript “$\operatorname*{isom}$” indicates that these conditions always lead to well-posed problems..
  1. 1.
    Condition $\operatorname*{IC}_{s}^{\operatorname*{Green}}$ is
    1. (a)
      for $s\in\mathbb{C}^{\bullet}_{\geq 0}$ given by
      $\lim_{r\rightarrow\infty}r^{(n-1)/2}\mathopen{}\mathclose{{}\left(\partial_{r}{u}\mathopen{}\mathclose{{}\left(r\xi}\right)+s{u}\mathopen{}\mathclose{{}\left(r\xi}\right)}\right)=0\quad\text{uniformly in }\xi\in S_{1},$ (3.3)
    2. (b)
      for $s=0$ and
      1. i.
        $n\geq 3$ given by
        $\mathopen{}\mathclose{{}\left|u\mathopen{}\mathclose{{}\left(r\xi}\right)}\right|\leq Cr^{2-n}\quad\text{as }r\rightarrow\infty\text{ uniformly in }\xi\in S_{1},$
      2. ii.
        $n=2$ given by
        $\exists b\in\mathbb{R}\qquad u\mathopen{}\mathclose{{}\left(r\xi}\right)-b\log r\leq Cr^{-1}\quad\text{as }r\rightarrow\infty\text{ uniformly in }\xi\in S_{1}.$ (3.4)
  2. 2.
    Condition $\operatorname*{IC}_{s}^{\operatorname*{isom}}$ is the same as condition $\operatorname*{IC}_{s}^{\operatorname*{Green}}$ for $s\in\mathbb{C}_{\geq 0}$ and $\mathopen{}\mathclose{{}\left(s,n}\right)\neq\mathopen{}\mathclose{{}\left(0,2}\right)$ while for $s=0$ and $n=2$, condition $\operatorname*{IC}_{0}^{\operatorname*{isom}}$ is given by
    $\mathopen{}\mathclose{{}\left|u\mathopen{}\mathclose{{}\left(r\xi}\right)}\right|\leq C\quad\text{as }r\rightarrow\infty\text{ uniformly in }\xi\subset S_{1}.$
In the following proposition we state the well-posedness of these exterior Dirichlet problems. We say that problem (3.1) is well posed for $\operatorname*{IC}_{s}$, if the corresponding solution operator $\operatorname{S_{D}}\in L(H^{1/2}(S_{R});H^{1}_{\mathrm{loc}}(B_{R}^{+}))$ that assigns each Dirichlet data $g\in H^{1/2}(S_{R})$ the unique solution $\operatorname{S_{D}}g\in H^{1}_{\mathrm{loc}}(B_{R}^{+})$ with $\operatorname{IC}_{s}$ to (3.1), is well-defined and continuous.
Proposition 3.2 (well-posedness).
  1. 1.
    Let $s\in\mathbb{C}^{\bullet}_{\geq 0}$.
    1. (a)
      Problem (3.1) with $\operatorname*{IC}_{s}^{\operatorname*{Green}}$ is well posed with a continuous solution operator denoted by $\operatorname*{S}_{\operatorname*{D}}\mathopen{}\mathclose{{}\left(s}\right)\in L(H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right);H_{\operatorname*{loc}}^{1}(B_{R}^{+}))$.
    2. (b)
      For any $g\in H^{1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)$ the solution $\operatorname*{S}_{\operatorname*{D}}\mathopen{}\mathclose{{}\left(s}\right)g$ also satisfies the stronger condition $\operatorname*{IC}_{s}^{\operatorname*{strong}}$. Hence (3.1) with $\operatorname*{IC}_{s}^{\operatorname*{strong}}$ is well posed with the solution operator $\operatorname*{S}_{\operatorname*{D}}\mathopen{}\mathclose{{}\left(s}\right)$.
  2. 2.
    Let $s=0$.
    1. (a)
      Then problem (3.1) with $\operatorname*{IC}_{0}^{\operatorname*{strong}}$ is well posed with a continuous solution operator denoted by ${\operatorname{S}_{\operatorname{D}}\mathopen{}\mathclose{{}\left(0}\right)}\in L(H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right);H_{\operatorname*{loc}}^{1}(B_{R}^{+}))$.
    2. (b)
      Let $s=0$ and $n=2$. If and only if $R\neq 1$, the problem (3.1) with $\operatorname*{IC}_{0}^{\operatorname*{Green}}$ is well posed with a continuous solution operator denoted by ${\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}\in L(H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right);H_{\operatorname*{loc}}^{1}(B_{R}^{+}))$.
Major part of the proof of Proposition 3.2 given in the appendix is a combination of results which are worked out in [13] in more generality.
Remark 3.3 (Lipschitz domains).
The results from Proposition 3.2 generalise to Dirichlet Helmholtz problems on the exterior of general bounded Lipschitz domains with the results in [13]. In this case, the condition $R\neq 1$ for the Laplace case in $n=2$ dimensions in Proposition 3.2.2b has to be replaced by the condition that the logarithmic capacity $\operatorname{Cap}_{\Gamma}$ defined in [13, p. 264] does not equal one, i.e., $\operatorname{Cap}_{\Gamma}\neq 1$.
3.2 The DtN operator for the Helmholtz problem on a sphere
The solution operators for the exterior Dirichlet Helmholtz problems from Proposition 3.2 define of the Dirichlet-to-Neumann operator. Recall that the normal derivative $\partial_{r}$ at $S_{R}$ is well defined for any solution to (3.1) by Remark 2.1. For any $s\in\mathbb{C}_{\geq 0}$, the operator $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right):H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)\rightarrow H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ is given by
$\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g:=\partial_{r}\operatorname*{S}\nolimits_{\operatorname*{D}}\mathopen{}\mathclose{{}\left(s}\right)g\qquad\text{for all }g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right).$
In the special case $s=0$ and $R\neq 1$ in $n=2$ dimensions, we additionally set
$\operatorname{DtN}^{\operatorname{log}}g:=\partial_{r}{\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}g\qquad\text{for all }g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right).$
For functions $v\in H_{\operatorname*{loc}}^{1}\mathopen{}\mathclose{{}\left(B_{R}^{+}}\right)$, we write $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)v$ for $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left.v}\right|_{S_{R}}}\right)$ and analogously for ${\operatorname*{DtN}\mathopen{}\mathclose{{}\left(0}\right)}$.
The Dirichlet-to-Neumann operators allows us to characterize the solutions to the exterior problems stated in Proposition 3.2 by its traces on the sphere $S_{R}$: For $s\in\mathbb{C}_{\geq 0}$, their definition implies for any Helmholtz-harmonic function $u\in H_{\operatorname*{loc}}^{1}(B_{R}^{+})$ (i.e., $-\Delta u+s^{2}u=0$ in $B_{R}^{+}$) that
  • condition $\operatorname*{IC}_{s}^{\operatorname*{strong}}$ holds if and only if its normal derivative satisfies
    $\begin{array}[c]{cl}\partial_{r}u=\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)u&\text{for }s\in\mathbb{C}_{\geq 0}\text{ and }u=\operatorname{S}_{\operatorname{D}}\mathopen{}\mathclose{{}\left(s}\right)(u|_{S_{R}}),\end{array}$
  • condition $\operatorname*{IC}_{0}^{\operatorname*{Green}}$ holds for $n=2,s=0$, and $R\neq 1$ if and only if
    $\partial_{r}u=\operatorname{DtN}^{\operatorname{log}}\mathopen{}\mathclose{{}\left(u|_{S_{R}}}\right)\quad\text{for }u={\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}\mathopen{}\mathclose{{}\left(u|_{S_{R}}}\right).$
The first main result states the key properties of the DtN operator including bounds on its real and imaginary parts. The symmetry and the boundedness, given here in terms of the natural wavenumber-indexed norms, are more standard and included with elementary proofs for completeness.
Theorem 3.4 (properties of $\operatorname*{DtN}$).
The operator $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\in L(H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right);H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right))$ depends continuously on the wavenumber $s\in\mathbb{C}^{\bullet}_{\geq 0}$ and admits a continuous extension to $\mathbb{C}_{\geq 0}$ with
$\lim_{\begin{subarray}{c}s\in\mathbb{C}^{\bullet}_{\geq 0}\\ s\rightarrow 0\end{subarray}}\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)={\operatorname*{DtN}\mathopen{}\mathclose{{}\left(0}\right)}.$ (3.5)
Any $g,h\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ satisfy
  1. (i)
    $\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{h}}\right\rangle_{S_{R}}=\mathopen{}\mathclose{{}\left\langle g,\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\overline{h}}\right\rangle_{S_{R}}=\mathopen{}\mathclose{{}\left\langle g,\overline{\operatorname*{DtN}\mathopen{}\mathclose{{}\left(\overline{s}}\right)h}}\right\rangle_{S_{R}},$
  2. (ii)
    $\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{h}}\right\rangle_{S_{R}}}\right|\leq C\mathopen{}\mathclose{{}\left(n}\right)\mathopen{}\mathclose{{}\left\|g}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right),s}\mathopen{}\mathclose{{}\left\|h}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right),s},$
  3. (iii)
    $\frac{n-2}{2}\mathopen{}\mathclose{{}\left\|g}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}^{2}\leq-\operatorname{Re}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{g}}\right\rangle_{S_{R}}}\right)\leq C\mathopen{}\mathclose{{}\left(n}\right)\mathopen{}\mathclose{{}\left\|g}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right),s}^{2},$
  4. (iv)
    $0<\mp\operatorname{Im}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{g}}\right\rangle_{S_{R}}}\right)\leq C(n)\|g\|_{H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right),s}^{2}\hfill\text{whenever }\pm\operatorname{Im}s>0\text{ and }g\neq 0$
The constant $C(n)>0$ exclusively depends on $R$ and the dimension $n\geq 2$.
The second main result is a Friedrichs-type inequality for functions with homogeneous DtN boundary conditions on $S_{R}$ in $n\geq 3$ dimensions. The proofs of Theorems 3.4 and 3.5 follow in Subsection 4.4 below based an explicit representation of the $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)$ operator derived in Section 4 in terms of its Fourier coefficients.
Theorem 3.5 (Friedrichs-type inequality).
Consider a bounded Lipschitz domain $\Omega\subset\mathbb{R}^{n}$ in $n\geq 3$ dimensions such that $S_{R}\subset\partial\Omega$ is relatively closed in $\partial\Omega$. Any $v\in H^{1}\mathopen{}\mathclose{{}\left(\Omega,S_{R}}\right)$ with $\partial_{r}v=\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)v$ on $S_{R}$ for some $s\in\mathbb{C}_{\geq 0}$ satisfies
$\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}\leq C_{\operatorname*{F}}\sqrt{\|\nabla v\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}+\mathopen{}\mathclose{{}\left\|\partial_{r}v}\right\|_{H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}^{2}}.$ (3.6)
The constant $C_{\operatorname*{F}}>0$ is independent of $s$ and exclusively depends on $\Omega$ and $R$.
A corollary for Helmholtz-harmonic functions on a Lipschitz domain $\Omega$ concludes this section. Recall the constants $C_{\omega,0},C_{\omega,1}$, and $C_{\mathrm{F}}$ from (2.3)–(2.4) and Theorem 3.5.
Corollary 3.6.
Consider a bounded Lipschitz domain $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ and let $C_{\operatorname*{F}}$ be as in Theorem 3.5. Any $v\in H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right)$ with the properties
$\begin{array}[c]{cl}\mathopen{}\mathclose{{}\left.\mathopen{}\mathclose{{}\left(-\Delta v+s^{2}v}\right)}\right|_{\omega}=0&\text{for some }\omega\subset\Omega\text{ such that }S_{R}\text{ is a relatively closed subset of }\partial\omega,\\ \partial_{r}v=\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)v&\text{on }S_{R}\text{ for some }s\in\mathbb{C}_{\geq 0}\end{array}$
satisfies
$\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}$ $\leq\max\mathopen{}\mathclose{{}\left\{1,\sqrt{1+C_{\omega,1}^{2}}C_{\operatorname*{F}}}\right\}\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right),s},$ (3.7)
$\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right),s}$ $\leq C_{\operatorname*{tr}}\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right),s}.$ (3.8)
In particular, if $\mathopen{}\mathclose{{}\left|s}\right|\leq 1/\sqrt{2\max\{1,C_{\omega,1}C_{\operatorname*{F}}\}}$, then
$\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}\leq 2C_{\operatorname{F}}\sqrt{1+C_{\omega,1}^{2}}\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}.$ (3.9)
The constant $C_{\operatorname*{tr}}>0$ exclusively depends on the trace constants $C_{\Omega,0}$, $C_{\omega,1}$ as well as on $C_{\operatorname*{F}}$.
Proof.
The estimate (3.7) is trivial for $\mathopen{}\mathclose{{}\left|s}\right|\geq 1$ by definition of the norm $\mathopen{}\mathclose{{}\left\|\bullet}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right),s}$. By assumption, the connected sphere $S_{R}\subset\partial\omega$ is relatively closed in $\partial\omega$ and thereby with a positive distance from $\partial\omega\setminus S_{R}$. For $\mathopen{}\mathclose{{}\left|s}\right|\leq 1$, Theorem 3.5 and (2.3) provide
$\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}$ $\leq C_{\operatorname*{F}}\sqrt{\mathopen{}\mathclose{{}\left(1+C_{\omega,1}^{2}}\right)\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}+C_{\omega,1}^{2}\mathopen{}\mathclose{{}\left\|\Delta v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}}.$ (3.10)
Since $\Delta v=s^{2}v$ in $\omega$ and $|s|^{4}\leq|s|^{2}$, (3.6) follows for $|s|\leq 1$ from
$\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}$ $\leq C_{\operatorname*{F}}\sqrt{\mathopen{}\mathclose{{}\left(1+C_{\omega,1}^{2}}\right)\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}+C_{\omega,1}^{2}\mathopen{}\mathclose{{}\left|s}\right|^{2}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}}\leq C_{\operatorname*{F}}\sqrt{1+C_{\omega,1}^{2}}\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right),s}.$ (3.11)
Let $\max\mathopen{}\mathclose{{}\left\{1,C_{\omega,1}C_{\operatorname*{F}}}\right\}\mathopen{}\mathclose{{}\left|s}\right|^{2}\leq 1/2$ (which implies $\mathopen{}\mathclose{{}\left|s}\right|\leq 1$) and observe from $\Delta v=s^{2}v$ in $\omega$ that
$\mathopen{}\mathclose{{}\left\|\Delta v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}=\mathopen{}\mathclose{{}\left|s}\right|^{2}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}(\omega)}\leq|s|^{2}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}.$ (3.12)
This and (3.10) establish the estimate (3.9) of the $L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)$ norm by
[…]
$\leq C_{\operatorname*{F}}\sqrt{1+C_{\omega,1}^{2}}\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}.$
It remains to prove the trace estimate in (3.8). An immediate consequence (known from [14, Cor. 3.2]) of a Young inequality and the trace estimate (2.4) reveals
$|s|\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}^{2}\leq\frac{C_{\Omega,0}^{2}}{2}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}+|s|^{2}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}}\right).$
Hence (3.8) results with $C_{\operatorname*{tr}}^{2}=3\mathopen{}\mathclose{{}\left(1+\max\mathopen{}\mathclose{{}\left\{1,\mathopen{}\mathclose{{}\left(1+C_{\omega,1}^{2}}\right)C_{\operatorname*{F}}^{2}}\right\}}\right)C_{\Omega,0}^{2}/2$ in
$\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right),s}^{2}\leq\frac{3C_{\Omega,0}^{2}}{2}\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}+\frac{C_{\Omega,0}^{2}}{2}|s|^{2}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}\leq C_{\operatorname*{tr}}^{2}\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right),s}^{2}.\qed$
Lemma 3.7 (special case for $n=2$).
Let $s=0$, $n=2$, and $R>1$. The difference of ${\operatorname*{DtN}\mathopen{}\mathclose{{}\left(0}\right)}$ for problem (3.1)–(3.2) and $\operatorname{DtN}^{\operatorname{log}}$ for problem (3.1) and (3.4) is constant, namely
${\operatorname*{DtN}\mathopen{}\mathclose{{}\left(0}\right)}g-\operatorname{DtN}^{\operatorname{log}}g=-\frac{1}{R\log R}\mathopen{}\mathclose{{}\left(\frac{1}{\mathopen{}\mathclose{{}\left|S_{R}}\right|}\int_{S_{R}}g}\right)\qquad\text{for all }g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right).$
The proof of Lemma 3.7 follows from an explicit Fourier expansion of the solutions ${\operatorname{S}_{\operatorname{D}}\mathopen{}\mathclose{{}\left(0}\right)}g$ and ${\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}g$ in Subsection 4.3.
4 Spectral expansion of the DtN operator
The $\operatorname*{DtN}$ operator has an explicit series expansion in $n$-dimensional polar coordinates in terms of Laplace-Beltrami eigenfunctions known for $s=\operatorname{i}\mathbb{R}$, e.g., from [2] for $n=2$ and [15, Eqn. (2.6.92)] for $n=3$ dimensions. An extension to general wavenumbers $s\in\mathbb{C}_{\geq 0}$ and dimensions $n\geq 2$ can be derived from [13, Chap. 9].
4.1 Spectral representation of the Helmholtz DtN
The spherical part of the DtN operator is expanded by an orthonormal basis of the eigenfunctions of the Laplace-Beltrami operator, whose properties are briefly recalled, e.g., from [17, Sec. 22.2]. The (negative) Laplace-Beltrami $-\Delta_{S_{1}}$ on the unit sphere $S_{1}\in\mathbb{R}^{n}$ is defined by
$-\Delta_{S_{1}}g:=-\mathopen{}\mathclose{{}\left.\mathopen{}\mathclose{{}\left(\Delta G}\right)}\right|_{S_{1}}$
with the extension of $g$ to $\mathbb{R}^{n}\backslash\mathopen{}\mathclose{{}\left\{0}\right\}$ by $G\mathopen{}\mathclose{{}\left(x}\right):=g\mathopen{}\mathclose{{}\left(x/\mathopen{}\mathclose{{}\left\|x}\right\|}\right)$. The eigenfunctions $Y_{m,j}$ (also called spherical harmonics) of the Laplace-Beltrami are given by
$-\Delta_{S_{1}}Y_{m,j}=\lambda_{m}Y_{m,j},\quad m\in\mathbb{N}_{0}\text{, }j\in J_{m},$ (4.1)
for some finite index set $J_{m}$ which corresponds to the multiplicity of the eigenvalue $\lambda_{m}$ for $m\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. We assume an ascending enumeration of the eigenvalues $\lambda_{m}$, i.e.,
$0=\lambda_{0}<\lambda_{1}<\ldots\overset{m\rightarrow\infty}{\longrightarrow}\infty,$
and that the eigenfunctions are chosen such that $\{Y_{m,j}\}_{j\in J_{m},m\in\mathbb{N}_{0}}$ forms a orthonormal basis of $L^{2}\mathopen{}\mathclose{{}\left(S_{1}}\right)$. Explicitly it holds [17, Thm. 22.1] that
$\lambda_{m}=m\mathopen{}\mathclose{{}\left(m+n-2}\right)\quad\text{with multiplicity }|J_{m}|=\frac{2m+n-2}{m+n-2}\binom{m+n-2}{n-2}.$ (4.2)
Next, we rewrite the exterior Helmholtz problem (3.1) in spherical coordinates $x=r\xi$, where $r:=\mathopen{}\mathclose{{}\left\|x}\right\|$ and $\xi:=x/r\in S_{1}$. Let $x=\psi\mathopen{}\mathclose{{}\left(r,\xi}\right)$ denote the associated transformation and write $\widehat{w}=w\circ\psi$ for a function $w$ in Cartesian coordinates. The Laplace operator in spherical coordinates takes the form
$\widehat{\Delta v}=\mathopen{}\mathclose{{}\left(\Delta v}\right)\circ\psi=\frac{1}{r^{n-1}}\partial_{r}\mathopen{}\mathclose{{}\left(r^{n-1}\partial_{r}\widehat{v}}\right)+\frac{1}{r^{2}}\Delta_{S_{1}}\widehat{v}.$
By expanding the solution $\widehat{u}=u\circ\psi$ to (3.1) by a Fourier series in terms of spherical harmonics
$\widehat{u}\mathopen{}\mathclose{{}\left(r,\xi}\right)=\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\widehat{u}_{m,j}\mathopen{}\mathclose{{}\left(r}\right)Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right),$ (4.3)
the Helmholtz equation (3.1) translates, to an ODE for the Fourier coefficients $\widehat{u}_{m,j}\mathopen{}\mathclose{{}\left(r}\right)$:
$-\frac{1}{r^{n-1}}\partial_{r}\mathopen{}\mathclose{{}\left(r^{n-1}\partial_{r}\widehat{u}_{m,j}\mathopen{}\mathclose{{}\left(r}\right)}\right)+\mathopen{}\mathclose{{}\left(\frac{m\mathopen{}\mathclose{{}\left(m+n-2}\right)}{r^{2}}+s^{2}}\right)\widehat{u}_{m,j}\mathopen{}\mathclose{{}\left(r}\right)=0\quad\text{for all }r>R.$ (4.4)
Let $I_{\mu}$ and $K_{\mu}$ denote the standard modified Bessel functions of order $\mu\in\mathbb{R}$ (e.g., from [4, §10]) and introduce the dimension-dependent parameter
$\nu=\frac{n-2}{2}\in\frac{1}{2}\mathbb{N}_{0}.$ (4.5)
The ODE (4.4) is closely related to the (modified) Bessel’s equation (see, e.g., [13, p. 278] for details). For any $s\in\mathbb{C}^{\bullet}_{\geq 0}$, a fundamental system is given by $f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(sr}\right)$ and $f_{m,\nu}^{(2)}\mathopen{}\mathclose{{}\left(sr}\right)$ with
$f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(z}\right):=\sqrt{\frac{\pi}{2}}\frac{K_{{}_{m+\nu}}\mathopen{}\mathclose{{}\left(z}\right)}{z^{\nu}}\qquad\text{and}\qquad f_{m,\nu}^{(2)}\mathopen{}\mathclose{{}\left(z}\right):=\frac{I_{{}_{m+\nu}}\mathopen{}\mathclose{{}\left(z}\right)}{\sqrt{2\pi}z^{\nu}}\qquad\text{for all }z\in\mathbb{C}^{\bullet}_{\geq 0}.$ (4.6)
Hence there exists coefficients $A_{m,j,\nu},B_{m,j,\nu}\in\mathbb{C}$ for the solution $\widehat{u}_{m,j}$ to (3.3) with
$\widehat{u}_{m,j}\mathopen{}\mathclose{{}\left(r}\right)=A_{m,j,\nu}\,f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(sr}\right)+B_{m,j,\nu}\,f_{m,\nu}^{(2)}\mathopen{}\mathclose{{}\left(sr}\right)\qquad\text{for all }r>R.$ (4.7)
Standard asymptotic expansions of modified Bessel functions [4, §10.40.1–4] for large argument provide
$\begin{array}[c]{ll}f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(z}\right)=\frac{\operatorname*{e}^{-z}}{z^{\nu+1/2}}\mathopen{}\mathclose{{}\left(1+\mathcal{O}\mathopen{}\mathclose{{}\left(z^{-1}}\right)}\right),&\frac{d}{dz}f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(z}\right)=-\frac{\operatorname*{e}^{-z}}{z^{\nu+1/2}}\mathopen{}\mathclose{{}\left(1+\mathcal{O}\mathopen{}\mathclose{{}\left(z^{-1}}\right)}\right),\\ f_{m,\nu}^{(2)}\mathopen{}\mathclose{{}\left(z}\right)=\frac{\operatorname*{e}^{z}}{z^{\nu+1/2}}\mathopen{}\mathclose{{}\left(1+\mathcal{O}\mathopen{}\mathclose{{}\left(z^{-1}}\right)}\right),&\frac{d}{dz}f_{m,\nu}^{(2)}\mathopen{}\mathclose{{}\left(z}\right)=\frac{\operatorname*{e}^{z}}{z^{\nu+1/2}}\mathopen{}\mathclose{{}\left(1+\mathcal{O}\mathopen{}\mathclose{{}\left(z^{-1}}\right)}\right)\end{array}$ (4.8)
with the Landau-$\mathcal{O}$ notation as $\mathbb{C}_{\geq 0}\ni z\rightarrow\infty$. Consequently, for any fixed $s\in\mathbb{C}^{\bullet}_{\geq 0}$ it holds
$r^{\nu+1/2}\mathopen{}\mathclose{{}\left(\frac{d}{dr}f_{m,\nu}^{\mathopen{}\mathclose{{}\left(j}\right)}\mathopen{}\mathclose{{}\left(sr}\right)+sf_{m,\nu}^{\mathopen{}\mathclose{{}\left(j}\right)}(sr)}\right)=\begin{cases}\mathcal{O}\mathopen{}\mathclose{{}\left(\frac{1}{r}}\right)&\text{if }j=1,\\ \mathcal{O}\mathopen{}\mathclose{{}\left(\operatorname*{e}{}^{sr}}\right)&\text{if }j=2\end{cases}\quad\text{as }r\rightarrow\infty$ (4.9)
and this implies that the function $u$ given by (4.3) satisfies condition $\operatorname*{IC}_{s}^{\operatorname*{Green}}$ if and only if $B_{m,j,\nu}=0$ in (4.7) for all $m\in\mathbb{N}_{0}$ and $j\in J_{m}$. Moreover, (4.9) and the left relation for $f_{m,\nu}^{\mathopen{}\mathclose{{}\left(1}\right)}$ in (4.8) reveal the implication
$\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}\implies\operatorname*{IC}\nolimits_{s}^{\operatorname*{strong}}.$ (4.10)
This and the Dirichlet boundary condition $\widehat{u}(R,\xi)=u\circ\psi(R,\xi)=g\circ\psi(R,\xi)$ at $r=R$ imply
$A_{m,j,\nu}=\frac{1}{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(sR}\right)}\widehat{g}_{m,j}$
for the Fourier coefficients $\widehat{g}_{m,j}$ of the Dirichlet data $g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ given by
$\widehat{g}(\xi)=\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\widehat{g}_{m,j}Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right).$ (4.11)
Hence, the solution operator $\operatorname*{S}_{\operatorname*{D}}\mathopen{}\mathclose{{}\left(s}\right)$ for the exterior Helmholtz problem (3.1) maps the Dirichlet data $\widehat{g}\equiv g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ in the form (4.11) to
$\widehat{u}\mathopen{}\mathclose{{}\left(r,\xi}\right)=\sum_{m\in\mathbb{N}_{0}}\frac{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(sr}\right)}{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(sR}\right)}\sum_{j\in J_{m}}\widehat{g}_{m,j}Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right).$ (4.12)
Its normal (radial) derivative results for all $s\in\mathbb{C}^{\bullet}_{\geq 0}$ in the $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)$ operator
$\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\,g:=\frac{1}{R}\sum_{m\in\mathbb{N}_{0}}z_{m,\nu}\mathopen{}\mathclose{{}\left(sR}\right)\sum_{j\in J_{m}}\widehat{g}_{m,j}Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right)\quad\text{with}\quad z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right):=s\frac{\frac{d}{ds}f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(s}\right)}{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(s}\right)}.$ (4.13)
4.2 Analysis of the spectral coefficients
The following theorem generalises and improves the results from [14, 15] for purely imaginary wavenumbers $s\in\operatorname{i}\mathbb{R}$ to general wavenumbers $s\in\mathbb{C}_{\geq 0}$ and spatial dimensions $n$.
Theorem 4.1 (characterisation of $z_{m,\nu}$).
For any $m\in\mathbb{N}_{0}$ and $\nu\in\tfrac{1}{2}\mathbb{N}_{0}$, the spectral coefficients $z_{m,\nu}$ are holomorphic in $\mathbb{C}^{\bullet}_{\geq 0}$ with continuous extension to $\mathbb{C}_{\geq 0}$ and
  1. (i)
    $\overline{z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)}=z_{m,\nu}(\overline{s})$ for all $s\in\mathbb{C}_{\geq 0}$.
Moreover, the following holds for all $s\in\mathbb{C}_{\geq 0}$ and $\mathopen{}\mathclose{{}\left(m,\nu}\right)\neq(0,0)$ for some constant $c_{2}>0$:
  1. (ii)
    $0\leq-\operatorname{Re}z_{0,0}\mathopen{}\mathclose{{}\left(s}\right)\leq\tfrac{1}{2}+\operatorname{Re}s$  and  $\nu+1/2\leq-\operatorname{Re}z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)\leq m+2\nu+\operatorname{Re}s$.
  2. (iii)
    If ±$\pm\operatorname{Im}s>0$, then   $0<\mp$$\operatorname{Im}$$z_{0,0}\mathopen{}\mathclose{{}\left(s}\right)\leq c_{2}+|\operatorname{Im}s|$  and  $0<\mp$$\operatorname{Im}$$z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)\leq\mathopen{}\mathclose{{}\left|{\operatorname{Im}}s}\right|$.
The following lemma prepares the proof of Theorem 4.1 below. The modulus of the Hankel function $H_{\mu}^{(1)}=J_{\mu}+\operatorname{i}Y_{\mu}$ of order $\mu\in\mathbb{R}$ reads
$M_{\mu}\mathopen{}\mathclose{{}\left(x}\right)=\sqrt{J_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)+Y_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)}\qquad\text{for all }x>0$ (4.14)
in the standard notation with the Bessel functions $J_{\mu}$ and $Y_{\mu}$ of first and second kind [4, §10].
Lemma 4.2 (estimates for Hankel modulus functions).
The modulus $M_{\mu}$ satisfies
$M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)$ $\leq-x\frac{d}{dx}\mathopen{}\mathclose{{}\left(M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)}\right)\leq 2\mu M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)$ for all $\text{for all }x$ $>0,\quad\text{for all }\mu\geq 1/2,$ (4.15)
$2\mu\,M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)$ $\leq-x\frac{d}{dx}\mathopen{}\mathclose{{}\left(M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)}\right)\leq M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)$ for all $\text{for all }x$ $>0,\quad\text{for all }\mu\in\mathopen{}\mathclose{{}\left[0,\frac{1}{2}}\right].$ (4.16)
(In particular, $M_{1/2}^{2}\mathopen{}\mathclose{{}\left(x}\right)=-x\frac{d}{dx}\mathopen{}\mathclose{{}\left(M_{1/2}^{2}\mathopen{}\mathclose{{}\left(x}\right)}\right)$.) The bounds in (4.15)–(4.16) hold with a strict inequality ($\leq$ replaced by $<$) whenever $\mu\not\in\{0,1/2\}$ and are sharp in the sense that
$\lim_{x\rightarrow 0}\mathopen{}\mathclose{{}\left(-x\frac{\frac{d}{dx}\mathopen{}\mathclose{{}\left(M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)}\right)}{M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)}}\right)=2\mu\qquad\text{and\qquad}\lim_{x\rightarrow\infty}\mathopen{}\mathclose{{}\left(-x\frac{\frac{d}{dx}\mathopen{}\mathclose{{}\left(M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)}\right)}{M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)}}\right)=1.$
Proof.
A holomorphic extension of $M_{\mu}^{2}$ in (4.14) to $\mathbb{C}_{>0}$ is known from Nicholson’s integral formula [19, §13.73] as
[…]
with the modified Bessel function $K_{0}$ of order 00. Since $\cosh$ and $\sinh$ are positive for positive arguments and increasing while $K_{0}\mathopen{}\mathclose{{}\left(x}\right)$ is decreasing for increasing $x>0$, it follows that $M_{\mu}^{2}(x)$ is decreasing in the real variable $x>0$ for fixed $\mu\geq 0$. A differentiation under the integral for positive $z\leftarrow x>0$ (both sides in (4.17) are analytic for $\operatorname{Re}z>0$) provides
[…]
Employ $\frac{d}{dt}K_{0}\mathopen{}\mathclose{{}\left(2x\sinh\mathopen{}\mathclose{{}\left(t}\right)}\right)=2x\cosh(t)\,K_{0}^{\prime}(2x\sinh(t))$ from the chain rule and
${g_{\mu}}\mathopen{}\mathclose{{}\left({t}}\right):={\tanh}\mathopen{}\mathclose{{}\left({t}}\right){\coth}\mathopen{}\mathclose{{}\left({2\mu\,t}}\right){\ =\frac{\sinh(t)}{\sinh(2\mu\,t)}\,\frac{\cosh(2\mu\,t)}{\cosh(t)}\ \qquad\text{for all }t\in(0,\infty)}$
to rewrite the last integral in the form
[…]
A straightforward calculation yields
$g_{\mu}^{\prime}\mathopen{}\mathclose{{}\left(t}\right)=\frac{d_{\mu}\mathopen{}\mathclose{{}\left(t}\right)}{\cosh^{2}\mathopen{}\mathclose{{}\left(t}\right)\sinh^{2}\mathopen{}\mathclose{{}\left(2\mu t}\right)}\quad\text{with\ }d_{\mu}\mathopen{}\mathclose{{}\left(t}\right):=w\mathopen{}\mathclose{{}\left(2\mu t}\right)-2\mu w\mathopen{}\mathclose{{}\left(t}\right)\quad\text{and}\quad w\mathopen{}\mathclose{{}\left(t}\right):=\sinh\mathopen{}\mathclose{{}\left(t}\right)\cosh\mathopen{}\mathclose{{}\left(t}\right)$
for all $t\in\mathbb{R}$. The derivative of $d_{\mu}$ satisfies
$d_{\mu}^{\prime}\mathopen{}\mathclose{{}\left(t}\right)=2\mu\mathopen{}\mathclose{{}\left(\cosh\mathopen{}\mathclose{{}\left(4\mu t}\right)-\cosh\mathopen{}\mathclose{{}\left(2t}\right)}\right)\qquad\text{for all }t\in\mathbb{R}.$
For $t>0$, the functions $\sinh\mathopen{}\mathclose{{}\left(t}\right)$ and $\cosh\mathopen{}\mathclose{{}\left(t}\right)$ are positive and increasing so these properties are inherited by $w\mathopen{}\mathclose{{}\left(t}\right)$. For $\mu>1/2$ we have $d_{\mu}^{\prime}\mathopen{}\mathclose{{}\left(t}\right)>0$ and $d_{\mu}\mathopen{}\mathclose{{}\left(0}\right)=0$ implies $d_{\mu}\mathopen{}\mathclose{{}\left(t}\right)>0$. This allows us to conclude that $g_{\mu}^{\prime}\mathopen{}\mathclose{{}\left(t}\right)$ is positive for $\mu>1/2$. An analogous reasoning verifies that $g_{\mu}^{\prime}\mathopen{}\mathclose{{}\left(t}\right)$ is negative for $\mu\in\mathopen{}\mathclose{{}\left]0,1/2}\right[$. Hence $g_{\mu}$ is monotonic and the limits $g_{\mu}\mathopen{}\mathclose{{}\left(t}\right)\rightarrow 1/\mathopen{}\mathclose{{}\left(2\mu}\right)$ as $t\rightarrow 0$ and $g_{\mu}\mathopen{}\mathclose{{}\left(t}\right)\rightarrow 1$ as $t\rightarrow\infty$ result in
$\min\mathopen{}\mathclose{{}\left\{1,1/\mathopen{}\mathclose{{}\left(2\mu}\right)}\right\}\leq g_{\mu}\mathopen{}\mathclose{{}\left(t}\right)\leq\max\mathopen{}\mathclose{{}\left\{1,1/\mathopen{}\mathclose{{}\left(2\mu}\right)}\right\}\quad\text{for all }t,\mu\in\mathopen{}\mathclose{{}\left(0,\infty}\right)$ (4.19)
with strict inequalities for $\mu\neq 1/2$. For $t>0$, the function $\sinh t>0$ is increasing while the function $K_{0}\mathopen{}\mathclose{{}\left(x}\right)$ is decreasing in $\mathopen{}\mathclose{{}\left(0,\infty}\right)$. Thus the term $\frac{d}{dt}\mathopen{}\mathclose{{}\left(K_{0}\mathopen{}\mathclose{{}\left(2x\sinh\mathopen{}\mathclose{{}\left(t}\right)}\right)}\right)<0$ is negative. Hence (4.18)–(4.19) reveal
$\max\mathopen{}\mathclose{{}\left\{1,2\mu}\right\}T_{\mu}\mathopen{}\mathclose{{}\left(x}\right)\leq x\frac{d}{dx}\mathopen{}\mathclose{{}\left(M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)}\right)\leq\min\mathopen{}\mathclose{{}\left\{1,2\mu}\right\}T_{\mu}\mathopen{}\mathclose{{}\left(x}\right)\quad\text{for all }x,\mu\in\mathopen{}\mathclose{{}\left(0,\infty}\right)$
(with strict inequality for $\mu\neq 1//2$) for
[…]
The series and the asymptotic expansion [4, §10.25.2, .31.2, .40.2] of $K_{0}$ result in
$\sin\mathopen{}\mathclose{{}\left(2\mu t}\right)K_{0}\mathopen{}\mathclose{{}\left(2x\sinh\mathopen{}\mathclose{{}\left(t}\right)}\right)\rightarrow 0\quad\text{as }t\rightarrow 0\text{ and as }t\rightarrow\infty.$
This leads to vanishing boundary terms in the integration by parts
[…]
The combination of (4.20)–(4.21) proves (4.15)–(4.16) for $x,\mu\in\mathopen{}\mathclose{{}\left(0,\infty}\right)$ (with strict inequality for $\mu\neq 1/2$). By continuity of $M_{\mu}^{2}=J_{\mu}^{2}+Y_{\mu}^{2}$ in $\mu$ ($J_{\mu}$ and $Y_{\mu}$ are entire in $\mu$ [4, §10.2(ii)]), (4.16) also holds for $\mu=0$. The exactness of the estimates in the limits $x\rightarrow 0$ and $x\rightarrow\infty$ follows from the series and the asymptotic expansions of $M_{\mu}^{2}\mathopen{}\mathclose{{}\left(x}\right)$, see, e.g., [4, §10.2, 10.8, and 10.18] for further details.∎
Proof of Theorem 4.1. Consider any $m\in\mathbb{N}_{0}$ and set $\mu:=m+\nu$ with $\nu$ from (4.5).
Step 1 departs with general remarks. The modified Bessel function $K_{m+\nu}=K_{\mu}$ has no zeroes [19, pp. 511–513] in $\mathbb{C}_{\geq 0}$ and is holomorphic [4, §10.25.ii] in $\mathbb{C}^{\bullet}_{\geq 0}$. Hence $z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)$ is well defined and holomorphic for $s\neq 0$. The product rule and the recurrence relations [4, §10.29.2] provide
$z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)=s\frac{K_{\mu}^{\prime}\mathopen{}\mathclose{{}\left(s}\right)}{K_{\mu}\mathopen{}\mathclose{{}\left(s}\right)}-\nu=m-s\frac{K_{\mu+1}(s)}{K_{\mu}(s)}.$ (4.22)
The behavior of $K_{\mu}$ for small $s\in\mathbb{C}^{\bullet}_{\geq 0}$ is known as
$K_{\mu}\mathopen{}\mathclose{{}\left(s}\right)=\mathopen{}\mathclose{{}\left\{\begin{array}[c]{lc}\frac{(\mu-1)!}{2}\mathopen{}\mathclose{{}\left(\frac{2}{s}}\right)^{\mu}+o\mathopen{}\mathclose{{}\left(s^{-\mu}}\right)&\text{if }\mu>0,\\ -\ln\mathopen{}\mathclose{{}\left(\frac{s}{2}}\right)-\gamma+o\mathopen{}\mathclose{{}\left(1}\right)&\text{if }\mu=0\end{array}}\right.$
(with Euler–Mascheroni constant $\gamma=0.5772\ldots$), e.g., from [4, §10.30.2, .31.2]. Hence
$\lim_{\begin{subarray}{c}s\in\mathbb{C}^{\bullet}_{\geq 0}\\ s\rightarrow 0\end{subarray}}s\frac{K_{\mu+1}\mathopen{}\mathclose{{}\left(s}\right)}{K_{\mu}\mathopen{}\mathclose{{}\left(s}\right)}=2\mu$
reveals that $z_{m,\nu}$ admits a continuous extension from $\mathbb{C}^{\bullet}_{\geq 0}$ to $\mathbb{C}_{\geq 0}$ by
$z_{m,\nu}\mathopen{}\mathclose{{}\left(0}\right):=\lim_{\begin{subarray}{c}s\in\mathbb{C}^{\bullet}_{\geq 0}\\ s\rightarrow 0\end{subarray}}z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)=m-2\mu=-\mathopen{}\mathclose{{}\left(m+2\nu}\right).$ (4.23)
For any real parameter $\mu\in\mathbb{R}$, the modified Bessel function $K_{\mu}(t)$ is real-valued for positive real arguments $t>0$ [4, §10.25.ii] and satisfies $\overline{K_{\mu}\mathopen{}\mathclose{{}\left(z}\right)}=K_{\mu}\mathopen{}\mathclose{{}\left(\overline{z}}\right)$ [4, §10.34.7] for all $z\in\mathbb{C}^{\bullet}_{\geq 0}$. Since complex conjugation commutes with differentiation, this and the definition of $z_{m,\nu}(s)$ prove (i).
Step 2 verifies (ii)–(iii) for purely imaginary wavenumbers $s=-\operatorname*{i}k$ with $k>0$. Recall the Bessel functions $J_{\mu}$ and $Y_{\mu}$ of first and second kind [4, §10.2]. The representation of $K_{\mu}(s)$ through Hankel functions $H_{\mu}^{(1)}(k)=J_{\mu}(k)+\operatorname{i}Y_{\mu}(k)$ [4, §10.27.8] is given by
$2K_{\mu}\mathopen{}\mathclose{{}\left(s}\right)=\pi\operatorname*{i}\nolimits^{\mu+1}{H_{\mu}^{(1)}}\mathopen{}\mathclose{{}\left(k}\right)\quad\text{with}\quad 2\frac{d}{ds}K_{\mu}(s)=\pi\operatorname*{i}\nolimits^{\mu+2}\frac{d}{dk}{H_{\mu}^{(1)}}\mathopen{}\mathclose{{}\left(k}\right)\qquad\text{for }s=-\operatorname{i}k.$
This and (4.22) result in
$z_{m,\nu}\mathopen{}\mathclose{{}\left(-\operatorname*{i}k}\right)=k\frac{\frac{d}{dk}H_{\mu}^{(1)}(k)}{H_{\mu}^{(1)}(k)}-\nu=w_{\mu}(k)-\nu\qquad\text{with}\quad w_{\mu}(k):=k\frac{\frac{d}{dk}H_{\mu}^{(1)}(k)}{H_{\mu}^{(1)}(k)}.$ (4.24)
The investigation of $w_{\mu}(k)$ departs with the decomposition into its real and imaginary parts known from [14, p. 1877] and is repeated here for convenience. First observe
$w_{\mu}(k)=k\frac{H_{\mu}^{(2)}(k)\frac{d}{dk}H_{\mu}^{(1)}(k)}{H_{\mu}^{(2)}(k)H_{\mu}^{(1)}(k)}=k\frac{J_{\mu}(k)J_{\mu}^{\prime}(k)+Y_{\mu}(k)Y_{\mu}^{\prime}(k)+\operatorname{i}\mathopen{}\mathclose{{}\left(J_{\mu}(k)Y_{\mu}^{\prime}(k)-Y_{\mu}(k)J_{\mu}^{\prime}(k)}\right)}{J_{\mu}^{2}(k)+Y_{\mu}^{2}(k)}$
by expanding (4.24) with $H_{\mu}^{(2)}(k)=J_{\mu}(k)-\operatorname{i}Y_{\mu}(k)$. The Wronskian identity $J_{\mu}(k)Y_{\mu}^{\prime}(k)-Y_{\mu}(k)J_{\mu}^{\prime}(k)=2/(\pi k)$ from [4, §10.5.2] is a consequence of the recurrence relations for Bessel functions. This reveals the characterisation [14, Eqn. (3.10)] of $w_{\mu}$, namely
[…]
Hence the real and imaginary parts of $z_{m,n}(-\operatorname{i}k)$ satisfy by (4.24) that
$-\operatorname{Re}z_{m,\nu}\mathopen{}\mathclose{{}\left(-\operatorname{i}k}\right)$ $=\nu-\operatorname{Re}w_{\mu}\mathopen{}\mathclose{{}\left(k}\right)=\nu-\frac{k}{2}\frac{\frac{d}{dk}\mathopen{}\mathclose{{}\left(M_{\mu}^{2}\mathopen{}\mathclose{{}\left(k}\right)}\right)}{M_{\mu}^{2}\mathopen{}\mathclose{{}\left(k}\right)},$ (4.25)
$\operatorname{Im}z_{m,\nu}\mathopen{}\mathclose{{}\left(-\operatorname{i}k}\right)$ $=\operatorname{Im}w_{\mu}\mathopen{}\mathclose{{}\left(k}\right)=\frac{2}{\pi M_{\mu}^{2}(k)}.$ (4.26)
Recall the assumption $k>0$ in this step. Lemma 4.2 controls the real part of $w_{\mu}(k)$ by
$\frac{1}{2}$ $\leq-\operatorname{Re}w_{\mu}(k)\leq\mu\qquad\text{for all }\mu\geq 1/2,$ (4.27)
00 0 $\leq-\operatorname{Re}w_{0}(k)\leq 1/2.$ (4.28)
Since $kM_{\mu}^{2}(k)$ is non-increasing for $\mu\geq 1/2$ and converges to $2/\pi$ as $k\rightarrow\infty$ [19, §13.74], one has $2\leq\pi kM_{\mu}^{2}(k)<\infty$ and
$20<\operatorname{Im}w_{\mu}(k)=\frac{2}{\pi M_{\mu}^{2}(k)}\leq k\qquad\text{for all }\mu\geq 1/$ 2 .
The situation is different for $\mu=0$ where $M_{0}^{2}(k)$ admits no lower bound for $k\rightarrow 0$ that is linear in $1/k$. However, the asymptotic expansion [19, §13.75] of $M_{0}^{2}(k)$ reveals the convergence $kM_{0}^{2}(k)\rightarrow 2/\pi$ as $k\rightarrow\infty$ with the lower bound (see [4, §10.18.iii])
$\frac{2}{\pi}\mathopen{}\mathclose{{}\left(1-\frac{1}{8k^{2}}}\right)\leq kM_{0}^{2}\mathopen{}\mathclose{{}\left(k}\right)\qquad\text{for all }k>0.$ (4.29)
Fix $k_{0}>2^{-3/2}$ and set $C_{0}:=\frac{k_{0}}{8k_{0}^{2}-1}$. This leads to
$\frac{2}{\pi M_{0}^{2}(k)}\leq k\frac{8k^{2}}{8k^{2}-1}=k+\frac{k}{8k^{2}-1}\leq k+C_{0}\qquad\text{for all }k>k_{0}.$ (4.30)
On the other hand, we have already seen that $M_{0}^{2}$ is non-increasing so that
$\sup_{0<k<k_{0}}\frac{2}{\pi M_{0}^{2}(k)}\leq\frac{2}{\pi M_{0}^{2}(k_{0})}=:C_{0}^{\prime}.$
For $c_{2}:=C_{0}+C_{0}^{\prime}$, this establishes
$00<\operatorname{Im}w_{0}(k)=\frac{2}{\pi M_{0}^{2}(k)}\leq c_{2}+k\qquad\text{ }\text{for all }k>$ 0 . (4.31)
The combination of (4.25)–(4.26) with (4.27) and (4.29) reads
$\nu+\frac{1}{2}\leq-\operatorname{Re}\,z_{m,\nu}\mathopen{}\mathclose{{}\left(-\operatorname{i}k}\right)\leq\nu+\mu\qquad\text{and}\qquad 0<{\operatorname{Im}}\,z_{m,\nu}\mathopen{}\mathclose{{}\left(-\operatorname{i}k}\right)\leq k$ (4.32)
for all $\mu\geq 1/2$, whereas (4.28) and (4.31) result for $\mu=0=m=\nu$ in
$0\leq-\operatorname{Re}\,z_{0,0}\mathopen{}\mathclose{{}\left(-\operatorname{i}k}\right)\leq\frac{1}{2}\qquad\text{and}\qquad 0<{\operatorname{Im}}\,z_{0,0}\mathopen{}\mathclose{{}\left(-\operatorname{i}k}\right)\leq c_{2}+k.$ (4.33)
Step 3 extends the analysis to general wavenumbers $s\in{\mathbb{C}}_{\geq 0}$. The behaviour of the modified Bessel functions for large $|s|$ and any fixed $\mu$ is known [4, §10.40.10] as
$K_{\mu}(s)=\sqrt{\frac{\pi}{2s}}\operatorname*{e}\nolimits^{-s}\mathopen{}\mathclose{{}\left(1+\frac{a_{1}(\mu)}{s}+R_{2}(\mu,s)}\right)$
where $a_{1}(\mu)=(4\mu^{2}-1)/8$ and the remainder $R_{2}(\mu,s)\leq C|s|^{-2}$ uniformly tends to zero for large $s\in\mathbb{C}_{\geq 0}$. Hence (4.22) reveals, for any $\mu=m+\nu\in\tfrac{1}{2}\mathbb{N}_{0}$, the asymptotic behaviour
$\lim_{\begin{subarray}{c}s\in\mathbb{C}_{\geq 0}\\ s\rightarrow\infty\end{subarray}}\mathopen{}\mathclose{{}\left(m-z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)-s}\right)=\lim_{\begin{subarray}{c}s\in\mathbb{C}_{\geq 0}\\ s\rightarrow\infty\end{subarray}}s\frac{K_{\mu+1}\mathopen{}\mathclose{{}\left(s}\right)-K_{\mu}(s)}{K_{\mu}\mathopen{}\mathclose{{}\left(s}\right)}=a_{1}(\mu+1)-a_{1}(\mu)=\mu+\frac{1}{2}.$
In other words, for any $\varepsilon>0$ there is some $r>0$ such that all $s\in\mathbb{C}_{\geq 0}$ with $|s|=r$ satisfy
$\nu+\frac{1}{2}-\varepsilon$ $\leq-\operatorname{Re}\,(z_{m,\nu}(s)+s)\leq-\operatorname{Re}\,z_{m,\nu}(s)\leq\nu+\frac{1}{2}+\operatorname{Re}\,s+\varepsilon,$
$-\varepsilon$ $\leq\mp\operatorname{Im}\,(z_{m,\nu}(s)+s)\leq\mp\operatorname{Im}\,z_{m,\nu}(s)\leq|\operatorname{Im}\,s|+\varepsilon\qquad\text{whenever }0\leq\pm\operatorname{Im}s.$
Since Step 2 controls the real part of $z_{m,\nu}$ on the imaginary axis (by (4.32)–(4.33) and $z_{m,\nu}(\overline{s})=\overline{z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)}$ from (i)), this leads for all $s\in\partial(B_{r}\cap\mathbb{C}_{\geq 0})$ to
$\nu+\frac{1}{2}-\varepsilon$ $\leq-\operatorname{Re}\,z_{m,\nu}(s)\leq\nu+\mu+\operatorname{Re}\,s+\varepsilon\qquad\text{if }1/2\leq\mu=m+\nu,$ (4.34)
00 0 $\leq-\operatorname{Re}\,z_{0,0}(s)\leq\frac{1}{2}+\operatorname{Re}\,s+\varepsilon.$ (4.35)
For the imaginary part, where $s\in\partial Q_{r}^{+}$ varies on the boundary of $Q_{r}^{+}:=\{s\in\overline{B_{r}}\cap\mathbb{C}_{\geq 0}\ :\ \operatorname{Im}s\geq 0\}$ in the first quadrant, it follows from (4.32)–(4.33) with (i) that
$-\varepsilon\leq-\operatorname{Im}z_{m,\nu}(s)\leq c_{n}+\operatorname{Im}s+\varepsilon$ (4.36)
with $c_{2}$ from (4.31) and $c_{n}:=0$ for all $n\geq 3$. It is well-known that the real and imaginary parts of the holomorphic function $z_{m,\nu}(s)$ (by Step 1) are harmonic. The maximum principle for harmonic functions [6, Thm. 2.2] reveals (4.34)–(4.35) for all $s\in B_{r}\cap\mathbb{C}_{\geq 0}$ and (4.36) for all $s\in Q_{r}^{+}$. Since this holds for all $\varepsilon>0$ and sufficiently large $r>0$, this concludes the proof of (ii) for the real part and leads for all $r>0$ to
$-\varepsilon\leq-\operatorname{Im}z_{m,\nu}(s)\leq c_{n}+\operatorname{Im}s+\varepsilon\qquad\text{for all }s\in Q_{r}^{+}.$
The strict sign $0<\mp\operatorname{Im}z_{m,\nu}(s)$ on the imaginary axis with $s\in\operatorname{i}\mathbb{R}\backslash\mathopen{}\mathclose{{}\left\{0}\right\}$ (by (4.32)–(4.33) and (i)) contradicts that $\operatorname{Im}z_{m,\nu}(s)\equiv 0$ vanishes in $Q_{r}^{+}$. Hence $0<-\operatorname{Im}z_{m,\nu}(s)$ also holds for all $s\in\mathrm{int}\,Q_{r}^{+}$ in the interior of $Q_{r}^{+}$ by the (strong) maximum principle. This and $\operatorname{Im}z_{m,\nu}(\overline{s})=-\operatorname{Im}z_{m,\nu}(s)$ from (i) conclude the proof of (iii).∎
4.3 Spectral representation of the Laplace DtN
The discussion in Section 4.1 reveals for $s=0$ that any $u\in H_{\operatorname*{loc}}^{1}(B_{R}^{+})$ with a series expansion of the form (4.3) solves the Laplace equation $-\Delta u=0$ if its Fourier coefficients $\widehat{u}_{m,j}$ given by (4.3) solve the ODE
$-\frac{1}{r^{n-1}}\partial_{r}\mathopen{}\mathclose{{}\left(r^{n-1}\partial_{r}\widehat{u}_{m,j}\mathopen{}\mathclose{{}\left(r}\right)}\right)+\frac{m\mathopen{}\mathclose{{}\left(m+n-2}\right)}{r^{2}}\widehat{u}_{m,j}\mathopen{}\mathclose{{}\left(r}\right)=0$ (4.37)
for all $r>R$ and $m\in\mathbb{N}_{0},j\in J_{m}$. A fundamental system for this ODE is given by
$f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(r}\right)=r^{-m-2\nu}\qquad\text{and}\qquad f_{m,\nu}^{(2)}\mathopen{}\mathclose{{}\left(r}\right):=\mathopen{}\mathclose{{}\left\{\begin{array}[c]{ll}r^{m}&\text{if }m+2\nu>0,\\ \log r&\text{if }m+2\nu=0\end{array}}\right.\qquad\text{for all }m\in\mathbb{N}_{0}.$
From this, we deduce that the behaviour of $f_{m,\nu}^{(2)}$ contradicts the condition $\operatorname*{IC}_{0}^{\operatorname*{strong}}$ while $f_{m,\nu}^{\mathopen{}\mathclose{{}\left(1}\right)}$ satisfies $\operatorname*{IC}_{0}^{\operatorname*{strong}}$. Hence, all coefficients $B_{m,\nu,j}$ in (4.7) must be zero. By matching the other coefficients with the series expansion of the Dirichlet data $g$ (with (4.11)), we obtain
$\widehat{u}\mathopen{}\mathclose{{}\left(r,\xi}\right)=\sum_{m\in\mathbb{N}_{0}}\frac{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(r}\right)}{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(R}\right)}\sum_{j\in J_{m}}\widehat{g}_{m,j}Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right)=\sum_{m\in\mathbb{N}_{0}}\mathopen{}\mathclose{{}\left(\frac{R}{r}}\right)^{m+2\nu}\sum_{j\in J_{m}}\widehat{g}_{m,j}Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right)\quad\text{for all }r>R,\xi\in S_{1}.$ (4.38)
Its normal (radial) derivative results in the $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)$ operator for the Laplacian
${\operatorname*{DtN}\mathopen{}\mathclose{{}\left(0}\right)}g=\frac{1}{R}\sum_{m\in\mathbb{N}_{0}}z_{m,\nu}^{\Delta}\sum_{j\in J_{m}}\widehat{g}_{m,j}Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right)\quad\text{with\quad}z_{m,\nu}\mathopen{}\mathclose{{}\left(0}\right):=R\frac{\frac{d}{dR}f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(R}\right)}{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(R}\right)}=-(m+2\nu).$
It follows from (4.23) in Step 1 of the proof of Theorem 4.1 that $z_{m,\nu}(0)$ is the limit $z_{m,\nu}(s)$ as $s\rightarrow 0$. Hence ${\operatorname*{DtN}\mathopen{}\mathclose{{}\left(0}\right)}$ is the continuous limit of $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)$ that converges as $s\rightarrow 0$ by Theorem 3.4. This proves the limit relation (3.5) in Theorem 3.4. Next we prove Lemma 3.7.
Proof of Lemma 3.7.
Consider problem (3.1) with $\operatorname*{IC}_{0}^{\operatorname*{Green}}$ for the Laplacian ($s=0$) and $n=2$, $R\neq 1$. Since $f_{0,0}^{(2)}$ (and not $f_{0,0}^{(1)}$) satisfies $\operatorname*{IC}_{0}^{\operatorname*{Green}}$, only the first term $m=0$ (with $J_{0}=\mathopen{}\mathclose{{}\left\{1}\right\}$) in the series representation (4.38) of the solution changes by the substitution
$\frac{f_{0,0}^{(1)}\mathopen{}\mathclose{{}\left(r}\right)}{f_{0,0}^{(1)}\mathopen{}\mathclose{{}\left(R}\right)}\leftarrow\frac{f_{0,0}^{(2)}\mathopen{}\mathclose{{}\left(r}\right)}{f_{0,0}^{(2)}\mathopen{}\mathclose{{}\left(R}\right)}=\frac{\log r}{\log R}.$
This leads to the representation
$\widehat{u}_{\mathrm{\log}}\mathopen{}\mathclose{{}\left(r,\xi}\right)=\frac{\log\mathopen{}\mathclose{{}\left(r}\right)}{\log\mathopen{}\mathclose{{}\left(R}\right)}\widehat{g}_{0,1}Y_{0,1}\mathopen{}\mathclose{{}\left(\xi}\right)+\sum_{m\in\mathbb{N}}\frac{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(r}\right)}{f_{m,\nu}^{(1)}\mathopen{}\mathclose{{}\left(R}\right)}\sum_{j\in J_{m}}\widehat{g}_{m,j}Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right)$ (4.39)
of the unique solution of Laplace problem (3.1) for $s=0$ with $\operatorname*{IC}_{0}^{\operatorname*{Green}}$. Since $Y_{0,1}(\xi)=Y_{0,1}$ is constant, the coefficient is explicitly given by the integral mean of $g$, namely
$\widehat{g}_{0,1}Y_{0,1}=\frac{Y_{0,1}^{2}\int_{S_{R}}g}{\|Y_{0,1}\|_{L^{2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}^{2}}=\frac{1}{|S_{R}|}\int_{S_{R}}g.$ (4.40)
Hence Proposition 3.2.2b reveals for any $R\neq 1$ that the function (4.39) is the unique solution ${\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}g$of the exterior Laplace problem with $\operatorname*{IC}_{0}^{\operatorname*{Green}}$. In the case $R=1$, the first summand in (4.39) is only well defined (as zero) provided $\int_{S_{1}}g=0$. Conversely, if $\int_{S_{1}}g\neq 0$ for $n=2$ and $R=1$, then (3.1) does not admit a solution with $\operatorname*{IC}_{0}^{\operatorname*{Green}}$.
It remains to consider the difference ${\operatorname{S}_{\operatorname{D}}\mathopen{}\mathclose{{}\left(0}\right)}g-{\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}g$ which can be written as
${\operatorname{S}_{\operatorname{D}}\mathopen{}\mathclose{{}\left(0}\right)}g-{\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}g=\widehat{g}_{0,1}Y_{0,1}-\frac{\log\mathopen{}\mathclose{{}\left(r}\right)}{\log\mathopen{}\mathclose{{}\left(R}\right)}\widehat{g}_{0,1}Y_{0,1}=-\frac{\log\frac{r}{R}}{\log R}\frac{1}{|S_{R}|}\int_{S_{R}}g.$
Hence, the difference of the $\operatorname*{DtN}$ operators satisfies
${\operatorname*{DtN}\mathopen{}\mathclose{{}\left(0}\right)}g-\operatorname{DtN}^{\operatorname{log}}g=-\frac{1}{R\log R}\frac{1}{|S_{R}|}\int_{S_{R}}g.$
This finishes the proof of Lemma 3.7. ∎
4.4 Proofs of Theorems 3.4 and 3.5
The sharp bounds in Theorem 4.1 and the explicit representation of the DtN operator (4.13) enable the proofs of Theorems 3.4 and 3.5 stated in Subsection 3.2.
Proof of Theorem 3.4.
This proof considers the unit disc $S_{1}$ ($R=1$) and the general case $R>0$ follows from the scaling of the norms and the arising constants then also depend on $R$. Any $g,h\in H^{1/2}(S_{1})$ with Fourier coefficients $\widehat{g}_{m,j},\widehat{h}_{m,j}$ from (4.11) with respect to the $L^{2}(S_{1})$ orthonormal basis $\{Y_{m,j}\}_{m\in\mathbb{N}_{0},j\in J_{m}}$ result in
$\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{h}}\right\rangle_{S_{1}}=\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}z_{m,\nu}(s)\,\widehat{g}_{m,j}\,\overline{\widehat{h}_{m,j}}=\mathopen{}\mathclose{{}\left\langle g,\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\overline{h}}\right\rangle.$ (4.41)
Since $z_{m,\nu}(s)=\overline{z_{m,\nu}(\overline{s})}$ by Theorem 4.1.i, this proves Theorem 3.4.i. Recall from the definition of the Fourier coefficients (4.11) and the equivalence of norms [11, Rem. 7.6] in $H^{1/2}(S_{1})$ that there exists a constant $C_{1}(n)>0$ that exclusively depends on $n$ such that
$\|g\|_{L^{2}(S_{1})}^{2}=\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\mathopen{}\mathclose{{}\left|\widehat{g}_{m,j}}\right|^{2}\quad\text{and}\quad\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\mathopen{}\mathclose{{}\left(1+m}\right)\mathopen{}\mathclose{{}\left|\widehat{g}_{m,j}}\right|^{2}\leq C_{1}\mathopen{}\mathclose{{}\left(n}\right)\mathopen{}\mathclose{{}\left\|g}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(S_{1}}\right)}^{2}.$ (4.42)
Theorem 4.1.ii–iii implies $\mathopen{}\mathclose{{}\left|\operatorname{Re}z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)}\right|\leq A+\operatorname{Re}s$ and $\mathopen{}\mathclose{{}\left|\operatorname{Im}z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)}\right|\leq B+\mathopen{}\mathclose{{}\left|\operatorname{Im}s}\right|$ for $A=\max\mathopen{}\mathclose{{}\left\{\frac{1}{2},m+2\nu}\right\}>0$ and $B=c_{2}>0$ such that
$\mathopen{}\mathclose{{}\left|z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)}\right|^{2}\leq\mathopen{}\mathclose{{}\left(A+\operatorname{Re}s}\right)^{2}+\mathopen{}\mathclose{{}\left(B+|\operatorname{Im}s|}\right)^{2}=\mathopen{}\mathclose{{}\left|A+\operatorname{Re}s+\operatorname*{i}\mathopen{}\mathclose{{}\left(B+|\operatorname*{Im}s|}\right)}\right|^{2}=\mathopen{}\mathclose{{}\left|A\pm\operatorname*{i}B+s}\right|^{2}.$
A triangle inequality results in
$\mathopen{}\mathclose{{}\left|z_{m,\nu}\mathopen{}\mathclose{{}\left(s}\right)}\right|$ $\leq\mathopen{}\mathclose{{}\left|A+\operatorname*{i}B+s}\right|\leq\mathopen{}\mathclose{{}\left|A}\right|+\mathopen{}\mathclose{{}\left|B}\right|+\mathopen{}\mathclose{{}\left|s}\right|$
$\leq m+2\nu+\frac{1}{2}+c_{2}+\mathopen{}\mathclose{{}\left|s}\right|\leq C_{2}\mathopen{}\mathclose{{}\left(n}\right)\mathopen{}\mathclose{{}\left(m+1}\right)+\mathopen{}\mathclose{{}\left|s}\right|$ (4.43)
for all $s\in\mathbb{C}_{\geq 0}$ with $C_{2}\mathopen{}\mathclose{{}\left(n}\right)=1+2\nu+c_{2}>0$ that only depends on $n=2\nu+2$.
Hence (4.41) and a Cauchy inequality verify for all $g,h\in H^{1/2}(S_{1})$ that
$|\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{h}}\right\rangle_{S_{1}}|$ $\leq\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\mathopen{}\mathclose{{}\left(C_{2}\mathopen{}\mathclose{{}\left(n}\right)\mathopen{}\mathclose{{}\left(1+m}\right)+\mathopen{}\mathclose{{}\left|s}\right|}\right)\,\mathopen{}\mathclose{{}\left|\widehat{g}_{m,j}}\right|\,\mathopen{}\mathclose{{}\left|\widehat{h}_{m,j}}\right|$
$\leq|s|\|g\|_{L^{2}(S_{1})}\mathopen{}\mathclose{{}\left\|h}\right\|_{L^{2}(S_{1})}+C_{1}(n)C_{2}\mathopen{}\mathclose{{}\left(n}\right)\|g\|_{H^{1/2}(S_{1})}\mathopen{}\mathclose{{}\left\|h}\right\|_{H^{1/2}(S_{1})}.$
This and a Cauchy inequality lead to the boundedness of $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\in L(H^{1/2}(S_{1});H^{-1/2}(S_{1}))$ in the form of Theorem 3.4.i with $C(n)=\max\{1,C_{1}(n)C_{2}\mathopen{}\mathclose{{}\left(n}\right)\}$. The continuity of the DtN operator in the wavenumber $s\in\mathbb{C}_{\geq 0}$ follows from Lebesgue’s dominated convergence theorem for the absolutely convergent series (4.41) with spectral coefficients $z_{m,\nu}(s)$ dominated by (4.43). Set $c\mathopen{}\mathclose{{}\left(n}\right)=\tfrac{n-2}{2}$. Theorem 4.1 controls the real and imaginary parts of (4.41) for $g=h$ and results with (4.42) in
$c(n)\|g\|_{L^{2}(S_{1})}^{2}$ $\leq-\operatorname{Re}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{g}}\right\rangle_{S_{1}}}\right)\leq\operatorname{Re}s\|g\|_{L^{2}(S_{1})}^{2}+C(n)\|g\|_{H^{1/2}(S_{1})}^{2},$
00 0 $<\mp\operatorname{Im}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{g}}\right\rangle_{S_{1}}}\right)\leq C(n)|\operatorname{Im}s|\|g\|_{L^{2}(S_{1})}^{2}\qquad\text{if }g\neq 0$
for some constant $C(n)$ that exclusively depends on $n$. This is (ii)–(iii) in Theorem 3.4 and concludes the proof. ∎
Proof of Theorem 3.5.
Recall the Hilbert space $H^{1}(\Omega,S_{R})$ from (2.5) (for $\Gamma=S_{R}$).
Assume for a contradiction to (3.6) that there exists two sequences $(v_{k})_{k\in\mathbb{N}}\subset H^{1}(\Omega,S_{R})$ and $(s_{k})_{k\in\mathbb{N}}\subset\mathbb{C}_{\geq 0}$ with $\partial_{r}v_{k}=\operatorname{DtN}(s_{k})v_{k}$ on $S_{R}$ such that $\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=1$ for all $k\in\mathbb{N}$ and $\|\nabla v_{k}\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}+\mathopen{}\mathclose{{}\left\|\partial_{r}v_{k}}\right\|_{H^{-1/2}(S_{R})}\rightarrow 0$ as $k\rightarrow\infty$. Since the sequence $(v_{k})_{k\in\mathbb{N}}$ is bounded in $H^{1}(\Omega,S_{R})$, it admits a weakly convergent subsequence with weak limit $u\in H^{1}(\Omega,S_{R})$. A standard compactness argument with the compact embedding $H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right)\hookrightarrow L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)$ and the weak lower semicontinuity of norms reveal $\mathopen{}\mathclose{{}\left\|u}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=1$ and $\mathopen{}\mathclose{{}\left\|\nabla u}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=0=\mathopen{}\mathclose{{}\left\|\partial_{r}u}\right\|_{H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}$. Hence $u\in H^{1}(\Omega,S_{R})$ is constant on the (in particular connected) domain $\Omega$.
The traces of $v_{k}$ and the constant $u$ on $S_{R}$ have the Fourier expansions
$v_{k}|_{S_{R}}=\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\widehat{v}_{m,j}^{(k)}Y_{m,j}\qquad\text{and}\qquad u|_{S_{R}}=\widehat{u}_{0,0}Y_{0,0}\quad\text{for }\widehat{u}_{0,0}=\frac{u}{Y_{0,0}}=\frac{u}{|S_{R}|^{1/2}}$
with $\widehat{v}_{\bullet,\bullet}^{(k)}\in\mathbb{C}$ for all $k\in\mathbb{N}$ in terms of the $L^{2}(S_{R})$-orthonormal Laplace-Beltrami eigenfunctions $Y_{\bullet,\bullet}$ from (4.1). The DtN operator representation (4.13) and $\partial_{r}v_{k}=\operatorname{DtN}(s_{k})v_{k}$ result in
$\frac{z_{0,\nu}(s_{k}R)}{R}\widehat{v}_{0,0}^{(k)}\,\overline{\widehat{u}_{0,0}}=\mathopen{}\mathclose{{}\left\langle\operatorname{DtN}(s_{k})v_{k},\overline{u}}\right\rangle_{S_{R}}=\mathopen{}\mathclose{{}\left\langle\partial_{r}v_{k},\overline{u}}\right\rangle_{S_{R}}\to 0$
with the weak convergence $\partial_{r}v_{k}\rightharpoonup\partial_{r}u=0$ in $H^{-1/2}(S_{R})$ as $k\to\infty$ in the last step. This and $1\leq\nu+1/2\leq|z_{0,\nu}(s_{k}R)|$ for $n\geq 3$ from Theorem 4.1.ii establish
$|\widehat{v}_{0,0}^{(k)}||\widehat{u}_{0,0}|\leq|z_{0,\nu}(s_{k}R)||\widehat{v}_{0,0}^{(k)}||\widehat{u}_{0,0}|=R\,\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left\langle\operatorname{DtN}(s_{k})v_{k},\overline{u}}\right\rangle_{S_{R}}}\right|\to 0\qquad\text{as }k\to\infty.$
Thus either $\widehat{u}_{0,0}=0$ or $\widehat{v}_{0,0}^{(k)}\to 0$ as $k\to\infty$ hold. The weak convergence $v_{k}|_{S_{R}}\rightharpoonup u|_{S_{R}}$ in $H^{1/2}(S_{R})$ (by the continuity of the trace map) implies the convergence $\widehat{v}_{0,0}^{(k)}\to\widehat{u}_{0,0}$ in $\mathbb{C}$ as $k\to\infty$ and reveals $\widehat{u}_{0,0}=0$ in the second case. Hence $u=0$ follows in any case. This contradicts $\|u\|_{L^{2}(\Omega)}=1$ and concludes the proof of (3.6). ∎
Appendix A Proof of Proposition 3.2
The proof of Proposition 3.2 follows mainly from the theory in [13], where integral equation techniques are used. The key role is played by Green’s representation formula and we sketch the arguments in the following. The starting point is the exterior problem [13, Eqn. (7.31)] ( for $\mathcal{P}=-\Delta+s^{2}$, $f=0$, and $\Gamma=S_{R}$ therein) written as:
find $\text{find }u\in H_{\operatorname*{loc}}^{1}\mathopen{}\mathclose{{}\left(B_{R}^{+}}\right)\quad\text{such that\quad}\mathopen{}\mathclose{{}\left\{\begin{array}[c]{rl}-\Delta u+s^{2}u=0&\text{in }B_{R}^{+},\\ u=g&\text{on }S_{R},\\ \mathcal{M}u=0&.\end{array}}\right.$
Here $\mathcal{M}u=0$ is a condition for the behaviour at infinity which is defined in [13, (7.29)]. Theorems 8.9 and 9.6 in [13] translate this condition for our setting to
$u\text{ satisfies }\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}\iff\mathcal{M}u=0.$
If the exterior problem (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ has a solution $u$, then this solution satisfies Green’s representation theorem by [13, Thm. 7.15], namely
$u=\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)g-\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)\psi\qquad\text{in }B_{R}^{+}$
where $\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)$ and $\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)$ are the single and double layer potential operators for the operator $\mathopen{}\mathclose{{}\left(-\Delta+s^{2}}\right)$ (see, e.g., [13, Eqn. (7.1)] for details) and $\psi$ is a solution of
$S\psi=\mathopen{}\mathclose{{}\left(-\frac{1}{2}+T}\right)g\qquad\text{on }S_{R},$ (A.1)
with the abbreviations $S:=\frac{1}{2}\mathopen{}\mathclose{{}\left(\gamma^{+}\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)+\gamma^{-}\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)}\right)$ and $T:=\frac{1}{2}\mathopen{}\mathclose{{}\left(\gamma^{+}\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)+\gamma^{-}\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)}\right)$. Note that for any $g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ and $h\in H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$, $\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)g$ and $\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)h$ define functions in $\mathbb{R}^{n}\backslash S_{R}$ and $\gamma^{+}$ denotes the trace from the outer domain $B_{R}^{+}$, while $\gamma^{-}$ is the trace from the inner domain $B_{R}$. It is proved in [13, Theorem 6.11] that the mappings
$\operatorname*{DL}\mathopen{}\mathclose{{}\left(s}\right)$ $:H^{1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)\rightarrow H_{\operatorname*{loc}}^{1}\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}\backslash S_{R}}\right),$
$\operatorname*{SL}\mathopen{}\mathclose{{}\left(s}\right)$ $:H^{-1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)\rightarrow H_{\operatorname*{loc}}^{1}\mathopen{}\mathclose{{}\left(\mathbb{R}^{n}}\right),$
$S$ $:H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)\rightarrow H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right),$
$T$ $:H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)\rightarrow H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$
are continuous linear operators. In this way, well-posedness of (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ follows if $S$ in (A.1) has a bounded inverse. Theorem 7.6 in [13] implies that $S$ is a Fredholm operator of index 00 so that uniqueness implies bounded invertibility. Uniqueness follows for $n\geq 2$ and $s\in\mathbb{C}^{\bullet}_{\geq 0}$ from [13, Theorem 9.11], for $n\geq 3$ and $s=0$ from [13, Corollary 8.11] and for $n=2$, $s=0$, and $R\neq 1$ from [13, Theorem 8.16.(ii)]. Hence (3.1) with condition $\operatorname*{IC}\nolimits_{s}^{\operatorname*{Green}}$ is well-posed in these three cases (with continuous solution operator). This finishes the proof of Proposition 3.2.1a and 3.2.2b, whereas 3.2.1b directly follows from the implication (4.10).
It remains to prove Proposition 3.2.2a. Uniqueness of solutions is guaranteed by [13, Thm 8.10]. Existence and well-posedness follows from the expansion (4.38) of the solution
$\widehat{u}\mathopen{}\mathclose{{}\left(r,\xi}\right)=\sum_{m\in\mathbb{N}_{0}}\mathopen{}\mathclose{{}\left(\frac{R}{r}}\right)^{m+2\nu}\sum_{j\in J_{m}}\widehat{g}_{m,j}Y_{m,j}\mathopen{}\mathclose{{}\left(\xi}\right)\quad\text{with\quad}u\mathopen{}\mathclose{{}\left(r\xi}\right)=\widehat{u}\mathopen{}\mathclose{{}\left(r,\xi}\right)$ (A.2)
We prove that the mapping $g\rightarrow u$ induced by this series is a continuous map from $H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ to $H_{\operatorname*{loc}}^{1}\mathopen{}\mathclose{{}\left(B_{R}^{+}}\right)$. For this, we show that for any $d>R$ it holds
$\mathopen{}\mathclose{{}\left|\kern-1.00006pt\mathopen{}\mathclose{{}\left|\kern-1.00006pt\mathopen{}\mathclose{{}\left|u}\right|\kern-1.00006pt}\right|\kern-1.00006pt}\right|_{H^{1}\mathopen{}\mathclose{{}\left(B_{R}^{+}\cap B_{d}}\right)}:=\mathopen{}\mathclose{{}\left(\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\mathopen{}\mathclose{{}\left(F_{m,j}^{2}+G_{m,j}^{2}+H_{m,j}^{2}}\right)\mathopen{}\mathclose{{}\left|\widehat{g}_{m,j}}\right|^{2}}\right)^{1/2}<\infty$ (A.3)
with
[…]
These coefficients allow for the straightforward estimate:
$F_{m,j}^{2}$$\leq\mathopen{}\mathclose{{}\left\{\begin{array}[c]{cc}\log\frac{d}{R}&m+\nu=0,\\ \frac{R^{2\nu}}{2(m+\nu)}&m+\nu\geq 1,\end{array}}\right.,\quad G_{m,j}^{2}\leq\mathopen{}\mathclose{{}\left(m+2\nu}\right)^{2}\mathopen{}\mathclose{{}\left\{\begin{array}[c]{ll}\frac{R^{2(\nu+1)}}{2(m+\nu-2)}&m+\nu\geq 3,\\ R^{2\mathopen{}\mathclose{{}\left(\nu+1}\right)}\log\frac{d}{r}&m+\nu=2,\\ \frac{1}{2}\mathopen{}\mathclose{{}\left(d^{2}-R^{2}}\right)R^{2\nu}&m+\nu=1,\\ \frac{1}{4}\mathopen{}\mathclose{{}\left(d^{4}-R^{4}}\right)R^{2\nu-2}&m+\nu=0,\end{array}}\right.$
$H_{m,j}^{2}$ =$=\mathopen{}\mathclose{{}\left\{\begin{array}[c]{ll}0&m+\nu=0,\\ \frac{m^{2}R^{2\nu}}{2(m+\nu)}&m+\nu\geq 1.\end{array}}\right.$
Consequently,
$\max\mathopen{}\mathclose{{}\left\{F_{m,j}^{2},G_{m,j}^{2},H_{m,j}^{2}}\right\}\leq\mathopen{}\mathclose{{}\left(\frac{d}{R}}\right)^{4}R^{2\nu}\mathopen{}\mathclose{{}\left(1+R^{2}}\right)\mathopen{}\mathclose{{}\left(m+2\nu}\right)\mathopen{}\mathclose{{}\left(\nu+2}\right).$
The combination of the previous arguments reveals
$\mathopen{}\mathclose{{}\left|\kern-1.00006pt\mathopen{}\mathclose{{}\left|\kern-1.00006pt\mathopen{}\mathclose{{}\left|u}\right|\kern-1.00006pt}\right|\kern-1.00006pt}\right|_{H^{1}\mathopen{}\mathclose{{}\left(B_{R}^{+}\cap B_{d}}\right)}$ $\leq\sqrt{3}\mathopen{}\mathclose{{}\left(\frac{d}{R}}\right)^{2}R^{\nu}\sqrt{\mathopen{}\mathclose{{}\left(1+R^{2}}\right)\mathopen{}\mathclose{{}\left(\nu+2}\right)}\mathopen{}\mathclose{{}\left(\sum_{m\in\mathbb{N}_{0}}\sum_{j\in J_{m}}\mathopen{}\mathclose{{}\left(m+2\nu}\right)\mathopen{}\mathclose{{}\left|\widehat{g}_{m,j}}\right|^{2}}\right)^{1/2}$
$\leq\sqrt{3}\mathopen{}\mathclose{{}\left(\frac{d}{R}}\right)^{2}R^{\nu}\sqrt{\mathopen{}\mathclose{{}\left(1+R^{2}}\right)\mathopen{}\mathclose{{}\left(\nu+2}\right)}\mathopen{}\mathclose{{}\left(1+2\nu}\right)\mathopen{}\mathclose{{}\left\|g}\right\|_{H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}.$
This implies for any $g\in H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ that the series (A.2) defines a continuous map to a function $u\in H_{\operatorname*{loc}}^{1}\mathopen{}\mathclose{{}\left(B_{R}^{+}}\right)$ and concludes the proof of existence and well-posedness in Proposition 3.2.2a.∎